OpenStudy (anonymous):

x^2+4x-5/5x^2-8x+3 multiplied by 20x-12/x^2-6x-55

OpenStudy (hoblos):

x^2+4x-5= (x+5)(x-1) 5x^2-8x+3= (5x-3)(x-1) x^2-6x-55= (x+5)(x-11) replace these and simplify

OpenStudy (anonymous):

How do you replace them?